# Dynamic neurons: A statistical physics approach for analyzing deep neural networks

### v1

@@ V Critical Phenomena
To make predictions about the scaling properties of general deep learning, we further study the fictitious fixed point from the example in the previous section, even though it is not physical. By considering the linear approximation of the RG transformation near the critical point, we can analyze the critical properties of the system.
$\begin{split}\vec{K}^{\prime}-\vec{K}^{*}&=\mathcal{R}(\vec{K})-\mathcal{R}(\vec{K}^{*})\\ &\approx\frac{d\mathcal{R}}{dK}\Large{|}_{K=K^{*}}(K-K^{*}),\end{split}$ (32)
@@
where $p=2$ denotes the ratio between the depth of the original and renormalized systems. The $y_{A}$, $y_{B}$, and $y_{C}$ determine the critical phenomena of the system. We can consider the continuum limit with respect to $p$. In other words, we can consider infinitesimal RG transformation from $M$ to $p^{-1}M$ where $p=1+\epsilon$.
@@
Considering $p^{y_{A}}|A-A^{*}|=A_{0}$ for some positive constant $A_{0}$, and substituting $p=(|A-A^{*}|/A_{0})^{-1/y_{A}}$, one can conclude that
@@
Now, considering that $r/p=r_{0}$ for some fixed distance $r_{0}$, and repeating the same argument, one can conclude that
@@
near the critical point.
These scaling laws are frequently observed in various examples of critical phenomena in statistical physics. Further research from the statistical physics perspective could provide deeper insights into these phenomena.
@@ VI Summary and Discussion
Our study introduced the dynamic neuron approach. In this approach, neurons are treated as dynamic variables of the system. We transformed the original Hamiltonian, where the degrees of freedom are synaptic weights and biases, into a new Hamiltonian with additional degrees of freedom—dynamic neurons.
The new Hamiltonian simplifies the interactions and reveals translational invariance. This symmetry, induced by the dynamic neuron approach, is approximate because the first and last layers are ignored. However, as the number of layers increases, the relative effects of the first and last layers diminish, and the symmetry becomes better preserved. This observation could be related to a fundamental question about DNNs: why do they perform so well when their depth is large? Our approach might offer insights into many fundamental questions about deep neural networks [59].
Additionally, the dynamic neuron approach highlights that the degrees of freedom—dynamic neurons, synaptic weights, and synaptic biases—interact primarily with their nearest neighbors. This interaction is intuitive, as neurons in a network communicate only with adjacent neurons to produce the final output. Consequently, the entire system can be understood by examining its individual components and their interactions.
We discussed the RG transformation and critical phenomena using a simple example of a DNN with dynamic neurons. Although it was a straightforward case, we demonstrated that well-known statistical approaches can be easily applied to DNNs using the dynamic neuron approach. RG transformations and critical phenomena are deeply studied topics in statistical physics with many physical implications. Finding connections between these implications and the properties of DNNs is very intriguing and promises rich outcomes.
@@ Acknowledgements.
Appendix A Spin Glass
In this appendix, we illustrate how the new Hamiltonian in the dynamic neuron approach replicates key characteristics of spin glass models, a class of statistical models renowned for their complex energy landscapes and disordered states. While there are several formulations of spin glass models, we focus on the Edwards-Anderson model, a canonical example widely used in statistical physics [60, 61]:
@@ Appendix A Spin Glass
where $s_{i}$ are spin degrees of freedom, which take values of $1$ or $-1$, and the coupling constants between spins $J_{ij}$ are typically assumed to follow independent probability distributions, such as Gaussian or Bernoulli distributions. $K_{i}$ is the coupling to magnetic fields which dependent on external magnetic field.

### v2

@@ V Critical Phenomena
To make predictions about the scaling properties of general deep learning, we further study the fictitious fixed point from the example in the previous section, even though it is not physical. By considering the linear approximation of the RG transformation near the critical point, we can analyze the critical properties of the system:
$\begin{split}\vec{K}^{\prime}-\vec{K}^{*}&=\mathcal{R}(\vec{K})-\mathcal{R}(\vec{K}^{*})\\ &\approx\frac{d\mathcal{R}}{dK}\bigg{|}_{K=K^{*}}(K-K^{*}),\end{split}$ (33)
@@
where $p=2$ denotes the ratio between the depth of the original and renormalized systems. [^14] In general, by taking the continuum limit of the system, the parameter $p$ can be extended to real numbers. This enables the definition of an infinitesimal RG transformation with parameter $p=1+\epsilon$, where $\epsilon$ is an infinitesimally small quantity. The values $y_{A}=1$, $y_{B}=1$, and $y_{C}=1$ determine the universality class of the system. The values may seem trivial here, but this is due to the simplicity of the system we are considering.
@@
Considering $p|A-A^{*}|=A_{0}$ for some positive constant $A_{0}$, and substituting $p=(|A-A^{*}|/A_{0})^{-1/y_{A}}$, one can conclude that
@@
Now, considering that $|m-n|/p=r_{0}$ for some fixed distance $r_{0}$, and repeating the same argument, one can conclude that
@@
near the critical point. [^16] Here, we define “distance” as the number of layers separating two neurons. In more complex deep neural network architectures, this concept may need a more refined definition. These scaling laws are frequently observed in various examples of critical phenomena in statistical physics. Further research from the statistical physics perspective could provide deeper insights into these phenomena.
Figure 7: Results of a Monte Carlo simulation for a simple neural network (red solid) and its dynamic neuron counterpart (blue dashed). We computed the time-averaged correlation function of neuron activations as a function of spatial separation, averaged over all positions. The two systems exhibit a close resemblance.
Equation (13) implies that, in the limit of large $\beta_{v}$, the DNA becomes equivalent to the DNN. To examine this correspondence, we performed a Monte Carlo simulation of the simplified neural network introduced in the previous section. The network depth was set to $10^{4}$, and we used $\beta=\beta_{v}=1$. As shown in Fig. 7, the time-averaged correlation functions of both systems exhibit a close resemblance. This observation suggests that, under suitable conditions, the DNA can serve as a reasonable theoretical approximation of the DNN. Owing to its analytical tractability, the DNA provides a useful perspective for further investigations into the structure and dynamics of deep neural networks.
@@ VI Summary and Discussion
Our study introduced the DNA. In this approach, neurons are treated as dynamic variables of the system. We transformed the original Hamiltonian, where the degrees of freedom are synaptic weights and biases, into a new Hamiltonian with additional degrees of freedom: dynamic neurons. The DNA with such a transformed system could be understood as a generalization of deep learning, where the limit of large $\beta_{v}$ corresponds to the typical DNN.
The new Hamiltonian simplifies the interactions and reveals translational invariance. This symmetry, induced by the DNA, is approximate because the first and last layers are ignored. However, as the number of layers increases, the relative effects of the first and last layers diminish, and the symmetry becomes better preserved. This observation could be related to a fundamental question about DNNs: why do they perform so well when their depth is large? Our approach might offer insights into many fundamental questions about DNNs [73].
Additionally, the DNA highlights that the degrees of freedom—dynamic neurons, synaptic weights, and synaptic biases—interact primarily with their nearest neighbors. This interaction is intuitive, as neurons in a network communicate only with adjacent neurons to produce the final output. Consequently, the entire system can be understood by examining its individual components and their interactions.
We discussed the RG transformation and critical phenomena using a simple example of a DNN with dynamic neurons. Although it was a straightforward case, we demonstrated that well-known statistical approaches can be easily applied to DNNs using the DNA. RG transformations and critical phenomena are deeply studied topics in statistical physics with many physical implications. Finding connections between these implications and the properties of DNNs is very intriguing and promises rich outcomes.
Although our study explains the statistical property of the specific example, it should be emphasized that the DNA can be applied to any DNN with arbitrary width, depth, and the number of data points. Many studies on DNNs concentrate on factors such as network size and the amount of training data. We anticipate that our DNA can become a useful tool to investigate those subjects.
In the present paper, given the limited theoretical development surrounding inference, we have focused on building a model that analytically explains the statistical properties of the training process. Addressing inference from a theoretical perspective remains challenging, as it involves understanding how neural networks generalize to unseen data sets. Future work should therefore develop theoretical tools for inference that would enable investigations into phenomena such as the neural scaling law [33, 34, 35, 36, 37, 38, 39, 40, 41, 42] and its relevance to our theory. These advances will help bridge the gap between the statistical properties of training and predictive performance in practical applications.
@@ Acknowledgements.
Appendix A Technical Application
This appendix delves into the technical implementation of the dynamic neuron system. Since the DNA modifies the Hamiltonian and introduces new degrees of freedom, referred to as “dynamic neurons,” it is crucial to analyze both the altered system dynamics and the behavior of these newly introduced components.
The new Hamiltonian described in Eq. (14) emerges from the dynamic neuron framework and governs the system’s behavior. Consequently, this introduces additional equations of motion for dynamic neurons and modifies the equations of motion for synaptic weights accordingly. To implement such a system, we can utilize the gradient descent method, as outlined in Eq. (3), in the following manner.
$\begin{split}&\Delta{W}_{ij}^{(m)}=-\eta\frac{\partial H_{\text{new}}}{\partial W_{ij}^{(m)}}\\ &\Delta{b}_{i}^{(m)}=-\eta\frac{\partial H_{\text{new}}}{\partial b_{i}^{(m)}}\\ &\Delta{h}_{i}^{(m)[l]}=-\eta\frac{\partial H_{\text{new}}}{\partial{h}_{i}^{(m)[l]}}\end{split}$ (41)
We need to implement changes for the dynamic neurons for each data represented by $l$. Under the lens of our theoretical framework, we anticipate that when $\beta_{v}$ is large enough the conventional DNN and the dynamic neuronal DNN will exhibit only minor differences in their statistical properties over the same training epochs.
However, it seems that constructing a DNN using the DNA is impractical at this point. In this framework, the number of dynamic neurons introduced is $LMN$, corresponding to the product of the number of data points, the width of the network and its depth. This implies that additional $LMN$ equations of motion must be solved to implement the system. Given the large datasets typically used in training, this approach appears computationally infeasible for practical applications.
Nonetheless, we remain open to the potential technical applications of this approach. Despite its current inefficiency, the development of better optimization strategies could enable the practical realization of the dynamic neuron framework. In many deep learning implementations, Eq. (3) are not solved simultaneously; instead, more efficient methods like backpropagation are commonly used for training. A similar method could be designed specifically for DNNs based on the DNA. The inherent homogeneity in the equations of motion, as highlighted in our main text, could serve as a basis for innovative optimization techniques. Moreover, since interactions in the DNA are restricted to nearest neighbors, the right-hand sides of Eq. (41) are expected to be simpler than those of Eq. (3). With further research into such optimization methods, we expect that the practical realization of the DNA could become feasible.
Appendix B Spin Glass
In this appendix, we illustrate how the new Hamiltonian in the DNA replicates key characteristics of spin glass models, a class of statistical models renowned for their complex energy landscapes and disordered states. While there are several formulations of spin glass models, we focus on the Edwards-Anderson model, a canonical example widely used in statistical physics [74, 75]:
@@ Appendix B Spin Glass
where $s_{i}$ are spin degrees of freedom, which take values of $1$ or $-1$, and the coupling constants between spins $J_{ij}$ are typically assumed to follow independent probability distributions, such as Gaussian or Bernoulli distributions. $K_{i}$ is the coupling to magnetic fields, which depends on an external magnetic field.